Jack usually mows his lawn in 5 hours. Mary can mow the same yard in 3 hours. how much time would it take for them to mow the lawn together? they could mow the lawn in ????
Answer by stanbon(57953)

Jack usually mows his lawn in 5 hours. Mary can mow the same yard in 3 hours. how much time would it take for them to mow the lawn together? they could mow the lawn in ????
-----------------
Jack rate: 1/5 job/hr
Mary rate: 1/3 job/hr
---
Together rate: 1/x job/hr
-----------------------------
Equation:
rate + rate = together rate
1/5 + 1/3 = 1/x
----
3x + 5x = 15
8x = 15
x = 15/8 = 1 7/8 hr
---
1 7/8 hr = 1hr (7/8)60 min = 1hr 52.5 minutes (time together)
=========================
Cheers,
Stan H.
==================